Dempsey Railroad Co. is about to issue $306,000 of 9-year bonds paying an 12% interest rate, with interest payable semiannually. The discount rate for such securities is 10%.

How much can Dempsey expect to receive for the sale of these bonds? (Round answer to 0 decimal places, e.g. 2,525.)
Answer #1
Amount PV factor 5% Present value
Semiannual interest 18360 11.68959 214621
Principal 306000 0.41552 127149
Total 341770
Amount to be received for the sale of these bonds = $ 341770
Workings:
Semiannual interest 18360 =306000*12%/2
PV factor 5%:
Semiannual interest 11.68959 =(1-(1.05)^-18)/0.05
Principal 0.41552 =1/1.05^18